OpenStudy (anonymous):

The dot product can also be used in applications such as in computing weighted averages. The syllabus of a college course weights material as follows: Attendance - 15%, HW - 15%, Quizzes - 10%, Exams - 40%, Final Exam - 20%. A student has an attendance of 35/42, a homework average of 78/100, quiz average of 17/25, and an exam average of 88%. What will the student's final numerical grade be if he scores an 84% on the final.

OpenStudy (anonymous):

multply and add

OpenStudy (anonymous):

\[.15\times \frac{35}{42}+.15\times .75+.1\times .68+.4\times .88+.2\times .84\]

OpenStudy (anonymous):

I think I understand....

OpenStudy (anonymous):

multiply each weight by each score then add them up

OpenStudy (anonymous):

Oh ok right....and that uses the dot product! Now I get it!